# What Can Large Language Models Do for Sustainable Food?

### v1

@@ Appendix B Extended Related Work
Our work broadly pertains to the integration of LLMs with other tools for reasoning tasks. Ahmaditeshnizi et al. (2024) develop an agent called OptiMUS for formulating and solving (mixed integer) linear programming problems from their natural language descriptions. Their goal is to expand access to solvers to individuals who lack the expertise required to formulate and solve these problems. Ye et al. (2024) propose SatLM, which combines LLMs with SAT solvers to solve problems in arithmetic reasoning, logical reasoning, and symbolic reasoning. Pan et al. (2023) propose Logic-LM, which integrates LLMs with symbolic solvers to improve logical reasoning. Relative to these works, our setting is fundamentally different, where the underlying optimization problem is not fully specified due to involving subjective human preferences or other components that are difficult to formalize mathematically. We are specifically motivated by the food domain, in which human satisfaction must be considered (alongside other factors such as nutrition, cost, emissions, etc.), but is difficult to write down mathematically. Similar problems arise in settings such as education (e.g. curriculum design) and health (e.g. fitness regimen design). Our framework is also applicable to settings where it may take considerable effort to formally specify the optimization problem, e.g. computing climate impacts as well as cost, nutrition, etc. for each food item, but an LLM may be able to generate sufficiently accurate estimates. We also note past work in the natural language processing literature on combining integer programming with machine learning, e.g. Denis & Baldridge (2007), which used integer linear programming for coference resolution.
@@ E.1 Experimental Design
Figures 11 and 12 show an example output from o1-preview.
@@ E.2 Menu Design
Emissions Satisfied Visit Recommend Animal Lives Plant Based LLM Gen. (AO) LLM Gen. (AC)
@@
o1-preview RD 11.05 7.48 6.98 6.76 0.18 0.36 0.50 0.20
@@
Table 12: Table of all results for the menu design task. “Emissions” is in kg of CO2eq per kg of food item. “Animal Lives” is number of animals killed per kg of food item. “Satisfied”, “Visit”, and “Recommend” are on a ten point (1 to 10) scale. “Plant Based” is the fraction of choices that were plant based (either vegan or vegetarian). “LLM Gen. (AO)” is the fraction of available options in the final optimized menu that were LLM generated (as opposed to recipes from the original menu). “LLM Gen. (AC)” is the fraction of actual choices that were LLM generated (as opposed to recipes from the original menu).
@@
Figure 13: Average emissions for all arms. $n=552$ across arms. Lower is better. Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview RD”, o1-preview’s descriptions are replaced with a simple list of ingredients.
@@
Figure 14: Average satisfaction for all arms. $n=552$ across arms. Higher is better. Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview RD”, o1-preview’s descriptions are replaced with a simple list of ingredients.
@@
Figure 15: Average likelihood of visiting the hypothetical restaurant for all arms. $n=552$ across arms. Higher is better. Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview RD”, o1-preview’s descriptions are replaced with a simple list of ingredients.
@@
Figure 16: Average likelihood of recommending the hypothetical restaurant to a friend for all arms. $n=552$ across arms. Higher is better. Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview RD”, o1-preview’s descriptions are replaced with a simple list of ingredients.
@@
Figure 17: Average animal usage for all arms. $n=552$ across arms. Lower is better for animal welfare. Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview RD”, o1-preview’s descriptions are replaced with a simple list of ingredients.
@@
Figure 18: Average emissions for all arms, excluding vegan and vegetarian participants. $n=514$ across arms. Lower is better. Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview RD”, o1-preview’s descriptions are replaced with a simple list of ingredients.
@@
Figure 19: Average satisfaction for all arms, excluding vegan and vegetarian participants. $n=514$ across arms. Higher is better. Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview RD”, o1-preview’s descriptions are replaced with a simple list of ingredients.
@@
Figure 20: Average likelihood of visiting the hypothetical restaurant for all arms, excluding vegan and vegetarian participants. $n=514$ across arms. Higher is better. Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview RD”, o1-preview’s descriptions are replaced with a simple list of ingredients.
@@
Figure 21: Average likelihood of recommending the hypothetical restaurant to a friend for all arms, excluding vegan and vegetarian participants. $n=514$ across arms. Higher is better. Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview RD”, o1-preview’s descriptions are replaced with a simple list of ingredients.
@@
Figure 22: Average animal usage for all arms, excluding vegan and vegetarian participants. $n=514$ across arms. Lower is better.Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview RD”, o1-preview’s descriptions are replaced with a simple list of ingredients.
@@ E.3 Sensory Profile Prediction
Accuracies in each quartile of the ground truth preference gap are shown in Tables 13141516.
@@ E.4 Recipe Preference Prediction
Tables 17 and 18 list supplementary analyses of performance on the recipe preference task. Table 19 shows performance on a smaller held-out set.

### v2

@@ Appendix B Extended Related Work
Our work broadly pertains to the integration of LLMs with other tools for reasoning tasks. Ahmaditeshnizi et al. (2024) develop an agent called OptiMUS for formulating and solving (mixed integer) linear programming problems from their natural language descriptions. Their goal is to expand access to solvers to individuals who lack the expertise required to formulate and solve these problems. Jiang et al. (2025) propose a five-element formulation for parameterizing optimization problems and fine-tune LLMs for generating five-element formulations and solver code from natural language descriptions of optimization problems. Ye et al. (2024) propose SatLM, which combines LLMs with SAT solvers to solve problems in arithmetic reasoning, logical reasoning, and symbolic reasoning. Pan et al. (2023) propose Logic-LM, which integrates LLMs with symbolic solvers to improve logical reasoning. Relative to these works, our setting is fundamentally different, where the underlying optimization problem is not fully specified due to involving subjective human preferences or other components that are difficult to formalize mathematically. We are specifically motivated by the food domain, in which human satisfaction must be considered (alongside other factors such as nutrition, cost, emissions, etc.), but is difficult to write down mathematically. Similar problems arise in settings such as education (e.g. curriculum design) and health (e.g. fitness regimen design). Our framework is also applicable to settings where it may take considerable effort to formally specify the optimization problem, e.g. computing climate impacts as well as cost, nutrition, etc. for each food item, but an LLM may be able to generate sufficiently accurate estimates. We also note past work in the natural language processing literature on combining integer programming with machine learning, e.g. Denis & Baldridge (2007), which used integer linear programming for coference resolution.
@@
Data-Driven Optimization
Our work also pertains to the literature on data-driven optimization. Previous work has proposed the “predict-then-optimize” framework in which a predictive model is used to define the parameters of an optimization problem (Elmachtoub & Grigas, 2022; Bertsimas & Kallus, 2020). We extend the predict-then-optimize framework in the setting of combinatorial optimization by incorporating a component where the elements of the ground set (e.g. recipes) are generated, and applying LLMs to both the generation and prediction steps.
@@ C.1 NECTAR Dataset
We also show some example pairs from this dataset.
Example Pair 1: Recipe 1 Corned Beef Dinner - Crock Pot Ingredients: corned beef brisket, onion, celery, potatoes, carrots, bay leaf, garlic clove, Worcestershire sauce, dry mustard, cabbage, caraway seed. Instructions: Trim brisket of all visible fat cut to fit 4 qt or larger crockpot if necessary. Place onion celery potatoes and carrots in bottom of crockpot lay brisket on top. Whisk together bouillon bay leaf garlic Worcestershire sauce and dry mustard. Pour over brisket cover pot. Cook on low setting for 8 to 10 hours adding cabbage wedges and caraway seed for the last hour of cooking. To serve discard cooking liquid slice meat onto hot serving plates accompany with the cooked potatoes carrots celery cabbage wedges and your favourite mustard.
Example Pair 1: Recipe 2 Crock Pot Corned Beef and Cabbage Ingredients: corned beef brisket, onions, cabbage, pepper, vinegar, sugar, water. Instructions: Combine ingredients in crock pot with cabbage on top. Cut meat to fit if necessary. Cover and cook on low 10-12 hours; high 6-7 hours or auto 6-8 hours.
Figure 4: Example pair 1 from the Food.com dataset.
Example Pair 2: Recipe 1 Oven-Fried Chicken Chimichangas Ingredients: salsa, ground cumin, dried oregano leaves, cheddar cheese, green onions, flour tortillas, margarine, cheddar cheese, green onion. Instructions: Mix chicken picante sauce or salsa cumin oregano cheese and onions. Place about 1/4 cup of the chicken mixture in the center of each tortilla. Fold opposite sides over filling. Roll up from bottom and place seam-side down on a baking sheet. Brush with melted margarine. Bake at 400°F for 25 minutes or until golden. Garnish with additional cheese and green onion and serve salsa on the side.
Example Pair 2: Recipe 2 Chicken Rice Casserole Ingredients: butter, flour, dried thyme, chicken broth, milk, salt, black pepper, cooked rice, parmesan cheese. Instructions: Melt 3 tablespoons of butter in a saucepan over medium heat whisk in flour and thyme and cook for 1 minute. Gradually stir in broth and milk stirring until thick and smooth. Stir in chicken and add salt and pepper; set aside. Grease a 1 1/2 quart shallow baking dish. Spread rice in prepared baking dish sprinkle with the peas and then pour creamed chicken mixture over. Dot with remaining tablespoon of butter and sprinkle with bread crumbs and cheese which you mix together. Bake at 400° until hot and bubbly about 20 to 25 minutes.
Figure 5: Example pair 2 from the Food.com dataset.
Example Pair 2: Recipe 1 Bow Tie Pasta With Broccoli and Broccoli Sauce Ingredients: broccoli, water, bow tie pasta, butter, salt. Instructions: Cut the broccoli buds into small florets leaving as short a stem as possible; set aside; cut the remaining broccoli stems into 1/2”̈ pieces; boil them covered in the 1 1/4 cups water lightly salted for 15 minutes or until very soft. Meanwhile boil the pasta in salted water until al dente; drain reserving 1/2 cup of the cooking water; return the pasta to the pot and stir until all moisture is evaporated. Place the reserved florets into a small pot and steam or boil 5 minutes retaining the bright green color and a bit of crunch. In a food processor puree the cooked broccoli and its cooking water until very smooth adding a bit of the reserved pasta water if needed to make a smooth sauce; add the butter and pulse until melted; add salt to taste. Pour warm sauce over the pasta and stir gently to combine; plate and sprinkle with drained florets.
Example Pair 2: Recipe 2 Lemon Garlic Pasta Ingredients: garlic, extra virgin olive oil, lemon juice, chicken flavor instant bouillon, pepper, spaghetti, angel hair pasta, parmesan cheese, parsley. Instructions: In large skillet cook garlic in olive oil until golden. Add lemon juice bouillon and pepper. Cook and stir until bouillon dissolves. In large bowl toss pasta garlic mixture cheese and parsley; serve immediately.
Figure 6: Example pair 3 from the Food.com dataset.
@@ E.1 Experimental Design
Figures 14 and 15 show an example output from o1-preview.
@@
Test Accuracy Complementarity Specificity Percent Time Saved
$t$-test 0.121 0.350 0.003 0.0002
Paired $t$-test 0.125 0.368 0.003 0.0003
Wilcoxon signed-rank test 0.078 0.284 0.005 0.0007
Linear mixed model with random effect for evaluator ID 0.075 0.993 0.001 0.0002
Permutation test 0.140 0.392 0.003 0.0003
OLS 0.1812 0.416 0.001 0.0006
Table 10: Sensitivity of $p$-values in the experimental design task to choice of statistical test. The paired $t$-test takes into account the pairing across products. The Wilcoxon signed-rank test is exact, for paired samples, nonparametric, and does not rely on normality assumptions. The permutation is also exact and does not rely on normality assumptions. In the OLS, we control for product ID, evaluator ID, generator ID as fixed effects.
@@ E.2 Menu Design
Emissions Satisfied Visit Recommend Animals PB LLM Gen. (AO) LLM Gen. (AC)
@@
o1-preview+IQP, RD 11.05 7.48 6.98 6.76 0.18 0.36 0.50 0.20
o1-preview+IQP, RP 9.44 7.84 7.63 7.33 0.12 0.55 0.36 0.39
o1-preview+IQP, RDi 9.75 8.16 7.96 7.66 0.13 0.50 0.47 0.34
o1-preview+IQP, RPDi 8.54 7.12 6.56 6.50 0.07 0.54 0.47 0.42
@@
Table 13: Table of all results for the menu design task. “Emissions” is in kg of CO2eq per kg of food item. “Animal Lives” is number of animals killed per kg of food item. “Satisfied”, “Visit”, and “Recommend” are on a ten point (1 to 10) scale. “PB” is the fraction of choices that were plant based (either vegan or vegetarian). “LLM Gen. (AO)” is the fraction of available options in the final optimized menu that were LLM generated (as opposed to recipes from the original menu). “LLM Gen. (AC)” is the fraction of actual choices that were LLM generated (as opposed to recipes from the original menu). In “o1-preview+IQP, RD”, o1-preview’s descriptions are replaced with a simple list of ingredients. In “o1-preview+IQP, RP”, the preferences component of the objective (the first term in Problem 2) is removed. In “o1-preview+IQP, RDi”, the diversity component of the objective is removed, i.e. $\lambda=0$ in Problem 2. In “o1-preview+IQP, RPDi”, both the preference and diversity components of the objective in Problem 2 are removed, i.e. it becomes a feasibility problem.
@@
Figure 16: Average emissions for all arms. $n=800$ across arms. Lower is better. Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview+IQP, Remove Descriptions”, o1-preview’s descriptions are replaced with a simple list of ingredients. In “o1-preview+IQP, Remove Preferences”, the preferences component of the objective (the first term in Problem 2) is removed. In “o1-preview+IQP, Remove Diversity”, the diversity component of the objective is removed, i.e. $\lambda=0$ in Problem 2. In “o1-preview+IQP, Remove Preferences and Diversity”, both the preference and diversity components of the objective in Problem 2 are removed, i.e. it becomes a feasibility problem.
@@
Figure 17: Average satisfaction for all arms. $n=800$ across arms. Higher is better. Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview+IQP, Remove Descriptions”, o1-preview’s descriptions are replaced with a simple list of ingredients. In “o1-preview+IQP, Remove Preferences”, the preferences component of the objective (the first term in Problem 2) is removed. In “o1-preview+IQP, Remove Diversity”, the diversity component of the objective is removed, i.e. $\lambda=0$ in Problem 2. In “o1-preview+IQP, Remove Preferences and Diversity”, both the preference and diversity components of the objective in Problem 2 are removed, i.e. it becomes a feasibility problem.
@@
Figure 18: Average likelihood of visiting the hypothetical restaurant for all arms. $n=800$ across arms. Higher is better. Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview+IQP, Remove Descriptions”, o1-preview’s descriptions are replaced with a simple list of ingredients. In “o1-preview+IQP, Remove Preferences”, the preferences component of the objective (the first term in Problem 2) is removed. In “o1-preview+IQP, Remove Diversity”, the diversity component of the objective is removed, i.e. $\lambda=0$ in Problem 2. In “o1-preview+IQP, Remove Preferences and Diversity”, both the preference and diversity components of the objective in Problem 2 are removed, i.e. it becomes a feasibility problem.
@@
Figure 19: Average likelihood of recommending the hypothetical restaurant to a friend for all arms. $n=800$ across arms. Higher is better. Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview+IQP, Remove Descriptions”, o1-preview’s descriptions are replaced with a simple list of ingredients. In “o1-preview+IQP, Remove Preferences”, the preferences component of the objective (the first term in Problem 2) is removed. In “o1-preview+IQP, Remove Diversity”, the diversity component of the objective is removed, i.e. $\lambda=0$ in Problem 2. In “o1-preview+IQP, Remove Preferences and Diversity”, both the preference and diversity components of the objective in Problem 2 are removed, i.e. it becomes a feasibility problem.
@@
Figure 20: Average animal usage for all arms. $n=800$ across arms. Lower is better for animal welfare. Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview+IQP, Remove Descriptions”, o1-preview’s descriptions are replaced with a simple list of ingredients. In “o1-preview+IQP, Remove Preferences”, the preferences component of the objective (the first term in Problem 2) is removed. In “o1-preview+IQP, Remove Diversity”, the diversity component of the objective is removed, i.e. $\lambda=0$ in Problem 2. In “o1-preview+IQP, Remove Preferences and Diversity”, both the preference and diversity components of the objective in Problem 2 are removed, i.e. it becomes a feasibility problem.
@@
Figure 21: Average emissions for all arms, excluding vegan and vegetarian participants. $n=740$ across arms. Lower is better. Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview+IQP, Remove Descriptions”, o1-preview’s descriptions are replaced with a simple list of ingredients. In “o1-preview+IQP, Remove Preferences”, the preferences component of the objective (the first term in Problem 2) is removed. In “o1-preview+IQP, Remove Diversity”, the diversity component of the objective is removed, i.e. $\lambda=0$ in Problem 2. In “o1-preview+IQP, Remove Preferences and Diversity”, both the preference and diversity components of the objective in Problem 2 are removed, i.e. it becomes a feasibility problem.
@@
Figure 22: Average satisfaction for all arms, excluding vegan and vegetarian participants. $n=740$ across arms. Higher is better. Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview+IQP, Remove Descriptions”, o1-preview’s descriptions are replaced with a simple list of ingredients. In “o1-preview+IQP, Remove Preferences”, the preferences component of the objective (the first term in Problem 2) is removed. In “o1-preview+IQP, Remove Diversity”, the diversity component of the objective is removed, i.e. $\lambda=0$ in Problem 2. In “o1-preview+IQP, Remove Preferences and Diversity”, both the preference and diversity components of the objective in Problem 2 are removed, i.e. it becomes a feasibility problem.
@@
Figure 23: Average likelihood of visiting the hypothetical restaurant for all arms, excluding vegan and vegetarian participants. $n=740$ across arms. Higher is better. Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview+IQP, Remove Descriptions”, o1-preview’s descriptions are replaced with a simple list of ingredients. In “o1-preview+IQP, Remove Preferences”, the preferences component of the objective (the first term in Problem 2) is removed. In “o1-preview+IQP, Remove Diversity”, the diversity component of the objective is removed, i.e. $\lambda=0$ in Problem 2. In “o1-preview+IQP, Remove Preferences and Diversity”, both the preference and diversity components of the objective in Problem 2 are removed, i.e. it becomes a feasibility problem.
@@
Figure 24: Average likelihood of recommending the hypothetical restaurant to a friend for all arms, excluding vegan and vegetarian participants. $n=740$ across arms. Higher is better. Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview+IQP, Remove Descriptions”, o1-preview’s descriptions are replaced with a simple list of ingredients. In “o1-preview+IQP, Remove Preferences”, the preferences component of the objective (the first term in Problem 2) is removed. In “o1-preview+IQP, Remove Diversity”, the diversity component of the objective is removed, i.e. $\lambda=0$ in Problem 2. In “o1-preview+IQP, Remove Preferences and Diversity”, both the preference and diversity components of the objective in Problem 2 are removed, i.e. it becomes a feasibility problem.
@@
Figure 25: Average animal usage for all arms, excluding vegan and vegetarian participants. $n=740$ across arms. Lower is better.Asterisks indicate a statistically significant difference ($t$-test with Bonferroni correction) compared to the original menu. Baselines are in green, ablations are in orange. In “o1-preview+IQP, Remove Descriptions”, o1-preview’s descriptions are replaced with a simple list of ingredients. In “o1-preview+IQP, Remove Preferences”, the preferences component of the objective (the first term in Problem 2) is removed. In “o1-preview+IQP, Remove Diversity”, the diversity component of the objective is removed, i.e. $\lambda=0$ in Problem 2. In “o1-preview+IQP, Remove Preferences and Diversity”, both the preference and diversity components of the objective in Problem 2 are removed, i.e. it becomes a feasibility problem.
@@ E.3 Sensory Profile Prediction
The performance of an expert food scientist specifically for the ‘Overall Satisfaction’ dimension is shown in Table 14. Accuracies in each quartile of the ground truth preference gap are shown in Tables 15161718.
Data Subset Expert Claude 3.5 Sonnet Gemini 1.5 Pro GPT-3.5 Turbo GPT-4o Llama o1-preview Nutr. Baseline
All 0.65 0.60 0.41 0.38 0.39 0.51 0.64 0.73
Quartile 1 0.62 0.43 0.29 0.43 0.33 0.43 0.43 0.71
Quartile 2 0.60 0.65 0.45 0.35 0.40 0.55 0.60 0.80
Quartile 3 0.77 0.64 0.59 0.41 0.64 0.59 0.68 0.55
Quartile 4 0.59 0.67 0.29 0.33 0.19 0.48 0.81 0.86
Table 14: Comparison of performance of the six LLMs on the sensory profile prediction task, specifically the “Overall Satisfaction” dimension, against an expert human food scientist. Statistically significant results (chi-squared test with Bonferroni correction) are in bold.
@@ E.4 Recipe Preference Prediction
Tables 19 and 20 list supplementary analyses of performance on the recipe preference task. Table 21 shows performance on a smaller held-out set.
@@ Appendix F Menus
  33. 33.
  34. 34.
  35. 35.
  36. 36.
Expert Chef
  1. 1.
    Crispy Miso Eggplant Ramen. Miso marinated eggplant in a crispy panko crust served with a shitake rich curry broth and hand pulled ramen noodles.
  2. 2.
    Pork Ribs. Pork Ribs smothered with Kentucky style BBQ sauce.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
    Star Anise Black Bean Tempeh Rolls. Black bean tempeh with a lacquered star anise hoisin sauce layered with pickled cucumber, shredded carrots, mint and maple toasted sesame seeds wrapped in rice paper.
  8. 8.
    Basque Celeriac and Gigante Bean Ragout. Celeriac, Gigante beans, nantes carrots, baby leeks, spring garlic, thyme, bay leaves and a lacto fermented tomato sauce creates an orchestra of flavors and textures.
  9. 9.
  10. 10.
    Beet and Tamari Infused “Tuna”. Watermelon infused with Tamari and beet juice, dehydrated and served on a beet horseradish puree with baby watercress, toasted tamari sunflower seeds and a yuzu dressing.
  11. 11.
  12. 12.
    Rustica Soyrizo Flatbread. Smoked paprika Soyrizo, toasted pinenut basil pesto slathered on extra virgin olive oil herb brushed house flatbread with confit sun dried tomato and garlic , spicy arugula.
  13. 13.
  14. 14.
  15. 15.
  16. 16.
    Beer Battered Shio Koji Marinated Enoki Mushroom. Shio koji marinated enoki mushroom, Okinawa sweet potatoes and Kabocha squash tempura with a fermented garlic dipping sauce.
  17. 17.
  18. 18.
  19. 19.
    Pasta e olio e ceci. cast iron sauteed garlic, chilli flakes and local pressed extra virgin olive oil tossed with al dente spaghetti and finished with pasta water and smashed ceci [garbanzo beans] then garnished with heaps of chopped Italian parsley and ground black pepper and a squeeze of Meyer lemon.
  20. 20.
    Supreme Pizza. Hand thrown sourdough pizza with fermented tomato basil sauce , wood grilled mushrooms and herbed cashew cheese.
  21. 21.
  22. 22.
  23. 23.
  24. 24.
  25. 25.
    Cheese Salad. Smoked cheddar, cheddar and sage stuffing.
  26. 26.
    Falafel with Tahini. Our signature recipe, served with a tahini dip.
  27. 27.
  28. 28.
    Vegan Meatballs. Vegan meatballs in a rich tomato and fennel sauce.
  29. 29.
  30. 30.
    Butternut Squash Salad. Roasted pomegranate molasses rubbed butternut squash tossed with pomegranate arils, toasted pumpkin seeds, herbed macadamia nut cheese with mushroom flat bread crumble.
  31. 31.
  32. 32.
    Pasta e lenticchie. Beluga lentil ragout slow cooked with Italian plum tomatoes ,tossed with bronze cut linguine and finished with smoky Shiitake bacon.
@@
Beef to Chicken
  1. 1.
  2. 2.
    Pork Ribs. Pork Ribs smothered with Kentucky style BBQ sauce.
  3. 3.
    Chicken Bourguignon. 6oz chicken burger topped with crispy bacon, red wine braised onions, Raclette cheese.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
    Slow Cooked Chicken in Chianti Sauce. Chicken and shallots slow cooked in rich Chianti red wine sauce with a cheesy dumpling.
  9. 9.
  10. 10.
  11. 11.
    Chicken and Venison Tagliatelle. A chicken breast, venison, red wine and beef dripping ragu with tagliatelle, fresh bufala mozzarella and crispy sage.
  12. 12.
  13. 13.
  14. 14.
  15. 15.
    Chicken Burger. 6oz chicken burger with baby gem lettuce and plum tomato in a chargrilled brioche bun with Dijon mayonnaise.
  16. 16.
    Beer Battered Fish with Chips. with frites and pea & mint puree.
  17. 17.
  18. 18.
    Baked Chicken. Baked chicken with red onion chutney, watercress, and Dijon mayonnaise.
  19. 19.
    Canelloni. pasta with béchamel, mozzarella and Gran Milano cheese.
  20. 20.
  21. 21.
  22. 22.
  23. 23.
  24. 24.
  25. 25.
    Cheese Salad. Smoked cheddar, cheddar and sage stuffing.
  26. 26.
    Falafel with Tahini. Our signature recipe, served with a tahini dip.
  27. 27.
  28. 28.
    Vegan Meatballs. Vegan meatballs in a rich tomato and fennel sauce.
  29. 29.
  30. 30.
  31. 31.
  32. 32.
  33. 33.
  34. 34.
  35. 35.
  36. 36.
@@
o1-preview, Remove Preferences
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
  11. 11.
  12. 12.
  13. 13.
  14. 14.
  15. 15.
    Canelloni. pasta with béchamel, mozzarella and Gran Milano cheese.
  16. 16.
    Beer Battered Fish with Chips. with frites and pea & mint puree.
  17. 17.
  18. 18.
  19. 19.
    Vegan Meatballs. Vegan meatballs in a rich tomato and fennel sauce.
  20. 20.
  21. 21.
    Pork Ribs. Pork Ribs smothered with Kentucky style BBQ sauce.
  22. 22.
  23. 23.
  24. 24.
    Falafel with Tahini. Our signature recipe, served with a tahini dip.
  25. 25.
  26. 26.
  27. 27.
  28. 28.
  29. 29.
  30. 30.
  31. 31.
  32. 32.
  33. 33.
  34. 34.
  35. 35.
  36. 36.
o1-preview, Remove Diversity
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
  11. 11.
  12. 12.
  13. 13.
  14. 14.
  15. 15.
  16. 16.
  17. 17.
  18. 18.
  19. 19.
  20. 20.
    Pork Ribs. Pork Ribs smothered with Kentucky style BBQ sauce.
  21. 21.
  22. 22.
  23. 23.
  24. 24.
  25. 25.
  26. 26.
    Beer Battered Fish with Chips. with frites and pea & mint puree.
  27. 27.
  28. 28.
  29. 29.
  30. 30.
  31. 31.
  32. 32.
    Falafel with Tahini. Our signature recipe, served with a tahini dip.
  33. 33.
  34. 34.
  35. 35.
  36. 36.
    Vegan Meatballs. Vegan meatballs in a rich tomato and fennel sauce.
o1-preview, Remove Preferences and Diversity
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
  11. 11.
  12. 12.
  13. 13.
  14. 14.
  15. 15.
  16. 16.
  17. 17.
  18. 18.
  19. 19.
    Beer Battered Fish with Chips. with frites and pea & mint puree.
  20. 20.
  21. 21.
  22. 22.
    Falafel with Tahini. Our signature recipe, served with a tahini dip.
  23. 23.
  24. 24.
  25. 25.
  26. 26.
  27. 27.
  28. 28.
    Pork Ribs. Pork Ribs smothered with Kentucky style BBQ sauce.
  29. 29.
    Canelloni. pasta with béchamel, mozzarella and Gran Milano cheese.
  30. 30.
  31. 31.
  32. 32.
  33. 33.
  34. 34.
  35. 35.
    Vegan Meatballs. Vegan meatballs in a rich tomato and fennel sauce.
  36. 36.